JeremyKoyels:

The perimeter of Dana's rectangular garden is 61 feet. The length of her garden is 5 feet more than twice the width. The perimeter of a rectangle can be found by adding the lengths of its four sides

mhanifa:

Do you know the perimeter formula?

Extrinix:

\(61=(2(w)+5)+(2(w)+5)+w+w\) So simplified would be: \(61=6w+10\) Now solve for the width, \(61=6w+10\) \(51=6w\) \(8.5=w\) Plug it in: \(61=(2(8.5)+5)+(2(8.5)+5)+8.5+8.5\) Which we can simplify into \(P=2l+2w\) \(61=2(22)+2(8.5)\) So, from this we get our length and width, \(l = 22\) \(w = 8.5\)